# Theorem-Carrying-Transaction: Runtime Certification to Ensure Safety for Smart Contract Transactions

### v1

@@ Abstract
Security bugs and trapdoors in smart contracts have been impacting the Ethereum community since its inception. Conceptually, the 1.45-million Ethereum’s contracts form a single “gigantic program” whose behaviors are determined by the complex reference-topology between the contracts. Can the Ethereum community be assured that this gigantic program conforms to its design-level safety properties, despite unforeseeable code-level intricacies? Static code verification is inadequate due to the program’s gigantic scale and high polymorphism. In this paper, we present a viable technological roadmap for the community toward this ambitious goal. Our technology, called Theorem-Carrying-Transaction (TCT), combines the benefits of concrete execution and symbolic proofs. Under the TCT protocol, every transaction carries a theorem that proves its adherence to the specified properties in the invoked contracts, and the runtime system checks the theorem before executing the transaction. Once a property is specified in a contract, it can be treated confidently as an unconditional guarantee made by the contract. As case studies, we demonstrate that TCT secures token contracts without foreseeing code-level intricacies like integer overflow and reentrancy. TCT is also successfully applied to a Uniswap codebase, showcasing a complex decentralized finance (DeFi) scenario. Our prototype incurs a negligible runtime overhead, two orders of magnitude lower than a state-of-the-art approach.
1 Introduction
Decentralized computing technologies have great potential to transform our society. However, a significant obstacle is the lack of robust assurance for smart contracts (i.e., application code executed on a blockchain platform). Numerous high-profile incidents have resulted in substantial financial losses due to unintentional vulnerabilities and deliberate logic trapdoors in smart contracts. These losses are irreversible due to the principle of code-is-law in decentralized computing [1]. For instance, the 2016 DAO attack led to a loss of 3.6 million ethers and a permanent hard fork of the Ethereum network [2]. The 2018 BeautyChain attack basically wiped out the entire asset base of the BeautyChain ecosystem [3]. Two attacks in 2022 exploited lending contracts to over-borrow ethers worth 11.7 million USD [4].
Vision. The community has developed techniques to detect code patterns leading to vulnerabilities, such as the integer-overflow and reentrancy patterns exploited by the BeautyChain and DAO attacks [2][3]. However, a more fundamental approach to trustworthiness is to ensure that an implementation conforms to its design-level safety properties, rather than to detect known patterns. Imagine traveling back to 2016 when the two patterns are still unknown. Suppose the designers of the ERC20 token standard [5] (which specifies the interface for fungible-token currencies) want to ensure that all ERC20 token contracts are secure against these unforeseen attacks, what can they do? It turns out that the property in Prop. 1 is sufficient to thwart both the DAO and BeautyChain attacks. This property asserts that every account balance (i.e., balances[x]) of this token contract should be non-negative, and the sum of all balances should equal the total supply of this currency (i.e., totalSupply). It simply states the designers’ intention, as opposed to enumerating all non-intentions as code-level anti-patterns. Section 3 will explain how this straightforward property defeats the unforeseen attacks.
[…]
Design-level properties are much more powerful than simple comparisons of concrete values (e.g.“x==y”). They may contain quantifiers, set/map operations, relations between encrypted variables, code-reflection properties (e.g., “no variable other than x can be modified”), etc. We envision a future in which contract designers and developers are empowered to express these powerful properties as the “higher law” to govern the contract code, and the platform ensures these properties so that the contracts are impervious to potential code bugs or trapdoors. We believe that a platform providing such high assurance will substantially bolster the public’s confidence on smart contract technology.
@@ 1 Introduction
Problem formulation. Smart contracts represent a unique programming paradigm. Conceptually, the whole community is co-developing a single gigantic “program”, as in Figure 1: everyone, even an attacker, can be a programmer and/or a user of this program; every contract is an OO object with a cryptographically strong identity called the “contract address”, i.e. an object reference. Any interaction between two objects is made by calling the target object via a reference. The platform (e.g. Ethereum VM, or EVM) sequentially (i.e. no concurrency) executes transactions, each of which is triggered by a user-invocation of a public method in an object. In 2022, the number of contracts on Ethereum’s mainnet had reached 1.45 million [6]. It is worth emphasizing that they are not 1.45 million individual programs, but a single program with 1.45 million objects interconnected!
@@
Figure 1: Smart contracts on Ethereum form a single gigantic “program” with about 1.45-million OO objects.
@@
For brevity, we refer to this gigantic program as $GP$. $GP$ has been self-evolving since 2015 and is highly polymorphic: (1) a transaction’s semantics is determined by the reference topology between the objects; (2) a transaction can create new contracts and references. The reference topology is a result of real-world business and trust relations, not implied by the static code. Therefore, static code alone is insufficient for reasoning about $GP$’s behaviors.
We use $P$ to denote the set of design-level properties co-defined by the whole community, which can be tens/hundreds of thousands of properties like Prop. 1. Hence, our technological goal can be formulated as: ensuring $GP$’s conformance to $P$. This formulation is very different from how a smart contract verification problem is typically formulated. Usually, a verification technique considers “the program” to be one or a few “contracts of interest”, denoted as $C$. $C$ is examined within a hypothetical environment, which is an enormous simplification of the huge $(GP-C)$ code in our formulation. It also assumes some hypothetical malicious users to invoke $C$ repeatedly. The verification serves as a helpful advisory to developers because it flags potential vulnerabilities, but the results are w.r.t. these vastly simplified assumptions. In contrast, our goal is a strong assurance – all properties in $P$ are indeed held true in reality, so everyone can trust them as strong contractual promises.
@@
Our technology. This is obviously an ambitious goal. Static verification is inadequate because of the gigantic scale and high polymorphism of $GP$. In this paper, we present a runtime certification approach and a roadmap for the Ethereum community to deploy it. The technology is named Theorem-Carrying-Transaction (TCT), inspired by Proof-Carrying-Authentication [7] and the concolic testing methodology [8]. In TCT, every transaction must carry a theorem that proves its adherence to the properties in every contract to be executed in the transaction. The runtime platform checks the validity of the theorem before allowing the transaction to be actually executed. The concolic nature of TCT means that it performs a symbolic verification on the concrete straight-line code trace of a transaction, which makes the corresponding theorem efficiently provable. Because all variables are symbolic, a proven theorem has the generality to cover all future transactions that follow the same code trace. This enables TCT to achieve near-zero amortized runtime overhead via a theorem reuse mechanism.
We will show that TCT achieves behavioral subtyping [9], a decades-long aspiration for secure OO programming. It mandates the properties of a base class to hold in all descendant classes. For example, if the designers specify Prop. 1 for the base class ERC20, they can rest assured that all future contracts derived from it will satisfy the property. Behavioral subtyping also gives the implementers a safety-certification mechanism. The fact that a contract is derived from a base class is essentially a certificate to prove to everyone that it meets the specified safety requirements.
@@
TCT’s trusted computing base (TCB) is entirely at the infrastructure level, i.e., the VM, the compiler, and the code verifier. It makes no assumption at the contract-code level, e.g., about the (static) code, the (dynamic) object-reference topology, or any party who claims that a contract has been formally verified. Of course, TCT shares the same caveat as all verification techniques – safety is only as high as the safety spec goes. We hope TCT motivates the community to shift focus from examining code-level intricacies to defining the design-level safety spec.
Prototype and highlights of results. We have built a prototype to demonstrate the TCT approach. In this prototype, a programmer uses the NatSpec notation of Solidity [10] to specify contract invariants and function post-conditions. TCT’s core module proves that these invariants and post-conditions will be satisfied under a given hypothesis (i.e., condition) w.r.t. a concrete EVM code trace (a.k.a. code path). This component consists of the techniques to obtain the code trace, to symbolically execute it and to convert the result into a straight-line code in a verification-amenable language (i.e., Boogie language [11] in our prototype). The hypothesis checking is performed on the Ethereum node by modifying the Go-Ethereum implementation (Geth) [12].
We demonstrate TCT using an ERC20 token contract that can be exploited by the integer overflow and reentrancy patterns. With intuitive hypotheses specified about the contract’s entry functions, our technology automatically proves the desired invariants, which thwart the attacks. To show its generality, the same technology is successfully applied to the transfer transactions of the top-5 real-world ERC20 tokens. Moreover, we showcase Uniswap-V2 as a highly complex real-world scenario. To the best of our knowledge, this is the first code-level verification for Uniswap’s essential properties – inverse proportionality for swap operations and proportionality for liquidity operations. Each transaction executes about 7000 $\sim$ 9000 EVM instructions across five different contracts. Indeed, we expose subtle conditions under which the properties fail to hold. They may lead to vulnerabilities if an investment strategy depends on Uniswap’s claimed properties.
@@
The runtime overhead of TCT is negligible. For the token contract and the Uniswap-V2 system, the overhead measurements are only 0.20% and 0.57% on Geth 1.13.0, respectively. The former is two orders of magnitude lower than a state-of-the-art approach, which inserts runtime checks into token contracts via code instrumentation [13].
2 Background
We briefly describe the techniques that inspire TCT.
PCC and PCA. Proof-Carrying Code (PCC) [14] is designed to allow a trusted party (e.g.an OS kernel) to safely accept a code module from an untrusted producer (e.g. a third-party driver). The novelty is that the producer bears the responsibility to prove the code’s conformance to the safety properties declared by the trusted party, who only needs to check the proof. Proof-Carrying Authentication (PCA) [7] is an authentication mechanism inspired by PCC. Real-world authentication policies, such as those involving “certificate authority”, “delegate”, the “speaksfor” relation, etc., are essentially statements about statements, so the authentication logic is higher-order, which is undecidable. PCA requires every request to carry a proof to convince the server that it is permitted by the authentication logic, so the server only needs to perform a decidable proof-checking task. PCA emphasizes its logic extensibility. Anyone can submit a proof of any theorem, e.g. “Charlie is a certificate authority”. Once the proof is confirmed, the theorem is accepted into the logic as a lemma, which can be used by the proofs of future theorems. Between PCC and PCA, the TCT concept is closer to the latter – a theorem attests to the safety of every runtime request rather than a code module.
Concolic testing. Another inspiration is from concolic testing, short for “concrete symbolic testing”, is a software testing technique that uses symbolic reasoning to analyze concrete code paths. It aims to systematically explore a program’s execution paths to increase code coverage and identify potential bugs. Concolic testing is particularly useful for finding complex and hard-to-detect software vulnerabilities, making it a valuable tool in software verification and security. For example, SAGE [15] is a concolic fuzz-testing technology that has found numerous bugs in Microsoft Windows and Office executables. The demonstrated scalability is inspiring for us because the code complexity of smart contracts is much lower than real-world binary executables.
3 Motivating Example
@@
To motivate our work, this section describes a token contract in Solidity that combines two vulnerabilities – the first is exploitable via integer overflow, similar to the one described in reference [16]; the second is exploitable via reentrancy, similar to the DAO attack [2]. In Figure 2, the middle portion shows the contract code MultiVulnToken. Smart contract execution is sequential and transactional. The require statement checks a condition: if it is false, the execution is reverted so that the EVM state remains the same as if the execution did not occur.
Figure 2: Attacks via integer overflow and reentrancy.
3.1 Vulnerabilities and attacks
@@
In Attack 1, the attacker issues a transaction by calling transferProxy (in transition A). Unexpectedly, arguments _value and _fee are set to $2^{255}$+1 and $2^{255}$. In the line require(balances[_from] >= _fee + _value), the two values are added, which yields result 1 due to integer overflow, so the require statement is satisfied, i.e., the execution is not reverted and the rest of the function also follows through. Address msg.sender is the attacker’s address, and address _to is also specified by the attacker. When the function returns, the balances of the two addresses become enormous (as transition B). This causes extreme inflation, so the attacker gets almost the entire asset of this token ecosystem.
@@
Attack 2 results in the attacker receiving 50 tokens in one of his accounts by clearing his other account that has only 5 tokens. It is achieved by making reentrant calls 10 times. The steps are as follows. First, the attacker creates a contract to call function clear in MultiVulnToken. Inside the function, msg.sender refers to the attacker’s contract. Therefore, the function call msg.sender.receiveNotification (transition C) calls into the attacker’s contract, who calls function clear again (transition D). This unexpected code sequence results in balances[_to] += bal executed 10 times.
@@
Consider from the perspective of a static code verifier. The behavior of clear is not fully defined because it calls into another contract using an address (i.emsg.sender). The verifier can perhaps err by conservatively flagging all such calls as “potentially unsafe”. However, calling contract via an address reference is very common in smart contracts. Imagine another call businessPartner.foo(), where businessPartner is a variable only writable by the contract owner. Treating it as an arbitrary address, thus flagging the code as “potentially unsafe”, is problematic. With its concolic nature, TCT is not subject to this dillema.
3.2 Theorem to prove safety of a transaction

### v2

# Theorem-Carrying Transactions: Runtime Verification to Ensure Interface Specifications for Smart Contract Safety

@@ Abstract.
Security bugs and trapdoors in smart contracts have been impacting the Ethereum community since its inception. Conceptually, the 1.45-million Ethereum’s contracts form a single “gigantic program” whose behaviors are determined by the complex compositions of contracts. Can programmers be assured that this gigantic program conforms to high-level safety specifications, despite unforeseeable code-level intricacies? Static code verification cannot be faithful to this gigantic program due to its scale and high polymorphism. In this paper, we present a viable approach to achieve this goal. Our technology, called Theorem-Carrying Transactions (TCT), combines the benefits of concrete execution and symbolic proofs. Under the TCT protocol, every transaction carries a theorem that proves its adherence to the specified properties in the invoked contracts, and the runtime system checks the theorem before executing the transaction. Once a theorem is proven, it will be reused for future transactions, so TCT’s runtime overhead is minimal. As case studies, we demonstrate that TCT secures token contracts without foreseeing code-level intricacies, such as integer overflow and reentrancy. TCT is also successfully applied to a Uniswap codebase, showcasing a complex decentralized finance (DeFi) scenario. Our evaluation shows a negligible runtime overhead, two orders of magnitude lower than a state-of-the-art approach for runtime checking of contract code safety.
copyright: none
1. Introduction
High assurance for smart contracts (i.e., application code executed on a blockchain platform like Ethereum) is crucial for the decentralized computing technology. The importance is evidenced by numerous high-profile incidents due to unintentional vulnerabilities and deliberate trapdoors in smart contracts (reentrancy_attacks, 25, 26, 28). These incidents caused substantial financial losses, which were irreversible due to the principle of code-is-law in decentralized computing (code-is-law, 5). To systematically tackle these issues, researchers have developed formal methods to ensure safety specifications for smart contracts.
Overview. In this paper, we present a technique named Theorem-Carrying Transactions (TCT). It offers a faithful and scalable verification approach for the Ethereum community to complement the assurance that existing (primarily static and compile-time) verification techniques can provide. Before explaining TCT’s assurance, we first present our “worldview” about smart contracts that differentiates TCT from existing techniques.
@@ 1. Introduction
Our “worldview” about smart contracts. Smart contracts represent a unique programming paradigm. Conceptually, the whole community is co-developing a single gigantic “program”, as in Figure 1: everyone, even an attacker, can be a programmer of it; every contract is an OO object with a cryptographically strong identity called the “contract address”, essentially an object reference. Any interaction between two objects is made by a cross-object method call via a reference. The platform (Ethereum VM, or EVM) sequentially (i.e., without concurrency) executes transactions, each triggered by a user-invocation of a public method in an object. In 2022, the number of contracts on Ethereum’s mainnet had reached 1.45 million (total-contracts, 1). We emphasize that they are not 1.45 million individual programs, but a single program with 1.45 million objects interconnected!
@@
Figure 1. Smart contracts on Ethereum form a single gigantic “program” with millions of OO objects.
@@
For brevity, we refer to this gigantic program, consisting of OO objects and the references between them, as $GP$. The $GP$ has been self-evolving since 2015 and is highly polymorphic: (1) the behavior of every cross-contract transaction is determined by the references held in the contracts; (2) a transaction can create new contracts and references. The reference topology is the result of real-world business and trust relationships, not implied by static code. In other words, the reference topology is the runtime data essential to the $GP$. When one talks about code-is-law of smart contracts, it is really the $GP$ one is referring to.
Our problem formulation. With this worldview, the problem we formulate is: how to ensure safety properties faithfully w.r.t. the entire $GP$? This is very different from how a smart contract verification problem is typically formulated. Usually, a verification technique considers “the program” to be one or a few “contracts of interest”, denoted $C$. $C$ is examined within a hypothetical environment of “external contracts”, which is an enormous simplification of the huge (GPC)(GP-C) code in reality. It also makes assumptions about the reference topology (whose ground truth is only available at runtime) and a hypothetical attacker to invoke $C$ with the attempt to violate safety specifications.
Static verification serves as a helpful advisory to programmers because it flags potential issues at compile time. However, with the environment code, reference topology and attacker all hypothetical, the verification is not faithful w.r.t. the $GP$. For soundness, it needs to err by assuming all possibilities of these hypothetical elements, which lead to false alarms, a well-known issue with static analysis.
Unlike static verification, a simple runtime assertion is actually faithful to the $GP$. Consider an assertion assert(x==y) in the source code. It is concretely checked when the line is reached, regardless of the aforementioned assumptions for static verification. TCT is designed to achieve the same faithfulness as runtime assertions. On the other hand, safety specifications that programmers want to ensure are more complex than checking concrete values. They usually specify contract invariants or method postconditions, which require symbolic reasoning. We refer to these as interface specifications as they specify properties at the method and contract levels rather than the individual statement level. TCT is designed to have the same symbolic-reasoning capability as static verification.
Interface specifications. The goal of TCT is to enable programmers to specify and ensure safety properties at the design level. This is different from detecting specific code-level patterns. For example, the 2016 DAO attack (reentrancy_attacks, 25) exploits an “integer overflow” pattern, and the 2018 BeautyChain attack (BEC_Attack, 28) exploits an “reentrancy” pattern. Researchers have developed techniques to detect these patterns.
Imagine traveling back to 2016 when the two patterns are still unknown. Suppose that the designers of the ERC20 token standard (ethereum-erc20, 33) (which specifies the interface for fungible-token currencies) want to ensure that all ERC20 token contracts are secure against these unforeseen attacks, what can they do? It turns out that the contract invariant in Property 1 is sufficient to thwart both the DAO and BeautyChain attacks. It asserts that every account balance (i.e., balances[x]) should be non-negative, and the sum of all balances should equal the total supply of this currency (i.e., totalSupply). It simply states the designers’ intention, as opposed to enumerating all non-intentions as code-level anti-patterns. Section 3 will explain how this straightforward property defeats the two attacks.
(1) $\begin{gathered}(\forall x:address|\ 0\leq this.balances[x]\leq this.totalSupply)\ \\ \wedge\ (\sum_{x}(this.balances[x])=this.totalSupply)\end{gathered}$
TCT ensures that: (1) a contract invariant holds before the contract is entered and after it exits; (2) a method postcondition holds at the method’s exit. As mentioned earlier, these specifications can be much more complex than checking basic values. They may contain quantifiers, set/map operations, relations between encrypted variables, code-reflection properties (e.g., “no variable except x can be modified”), real-number arithmetic with infinite precision, etc. All of these require symbolic reasoning about contract code, which is traditionally done by static verification.
@@
The TCT technology. TCT empowers symbolic verification of interface specifications to be scalable, efficient and faithful to the entire $GP$. It is inspired by Proof-Carrying Authentication (PCA, 2) and the concolic testing methodology (e.g., SAGE (godefroid2008automated, 9)). In TCT, every transaction must carry a theorem that proves its adherence to the properties in every contract to be executed in the transaction. The runtime platform checks the validity of the theorem before allowing the transaction to be executed. The concolic nature of TCT means that it performs a symbolic verification on the concrete straight-line code trace of a transaction, which makes the corresponding theorem efficiently provable. Because all variables are symbolic, a proven theorem has the generality to cover all future transactions that follow the same code trace. This enables TCT to achieve near-zero amortized runtime overhead via a theorem reuse mechanism.
The interface specifications are defined using the NatSpec notation of Solidity (NatSpec, 31). TCT’s verifier consists of the techniques to obtain a runtime EVM-code trace, to insert the specifications into the trace and to convert the trace into a straight-line code in a verification-amenable language (i.e., Boogie language (leino_boogie_2, 13) in our implementation). In addition to the verifier, TCT has its runtime mechanism, which is implemented in Go-Ethereum (Geth) (geth, 20).
@@
TCT’s trusted computing base (TCB) is entirely at the infrastructure level, i.e., the VM, the compiler, and the code verifier. It makes no assumptions at the contract-code level, e.g., about the (static) code, the (dynamic) object references, or any party who claims that a contract has been formally verified. Of course, TCT shares the same caveat as all verification techniques – safety is only as high as the safety specifications go. We hope TCT motivates the community to shift focus from examining code-level intricacies to defining safety specifications.
Results. To show the uniqueness of TCT, we begin with classical scenarios of ERC20 contracts. We show how TCT differs from existing techniques when tackling integer overflow and reentrancy attacks. TCT only requires programmers to define intuitive guard conditions for contracts’ entry methods, without examining code patterns of non-intentions. The TCT infrastructure ensures (1) these conditions logically imply the interface specifications; (2) these conditions are true at runtime. To show generality, our approach is applied on the top-5 real-world ERC20 tokens.
A crucial part of our evaluation is the case study of a highly complex real-world scenario – decentralized finance (DeFi). We showcase Uniswap-V2, which implements the core algorithm of Ethereum’s top DeFI application – Uniswap. To the best of our knowledge, this is the first faithful code-level verification for Uniswap’s essential properties – inverse proportionality for swap operations and proportionality for liquidity operations. We define them as interface specifications. Each transaction executes about 7000 $\sim$ 9000 EVM instructions across five different contracts. On one hand, the proven theorems explicate the precise cross-contract conditions that guard these properties. On the other hand, they expose subtle conditions under which the properties fail to hold. Accomplishing formal verification at this scale, we believe we are contributing a significant case study to the smart contract verification community.
@@
The runtime overhead of TCT is negligible. For the token contract and the Uniswap-V2 system, the overhead measurements are only 0.20% and 0.57% on Geth 1.13.0, respectively. The former is two orders of magnitude lower than a state-of-the-art approach, which inserts runtime checks into contracts via code instrumentation (li2020securing, 14).
2. Background
2.1. The Ethereum system and transactions
As shown in Figure 2, a user interacts with Ethereum via a transaction (TX) issuer (e.g., the MetaMask browser extension (metamask, 11)). Ethereum consists of a peer-to-peer network that runs a consensus protocol. Any machine can join the network to become a node. All nodes form the world-wide consensus community. The purpose of the protocol is to conceptually create a single VM (i.e., EVM) that is trustworthy if the compliant nodes constitute a majority in the community. This is achieved by replicating the EVM on every node.
The EVM is similar to the Java Virtual Machine (JVM). It runs bytecode, usually compiled from an object-oriented (OO) language like Solidity. However, there are a few differences: (1) a JVM runs many programs, each consisting of many objects, but an EVM runs a single program (i.e. the GP) consisting of all objects (i.e. contracts). It is important to note that the GP is not distributed across the nodes, but fully replicated on each and every node. (2) Executing a transaction on an EVM costs a certain amount of ether, referred to as gas. (3) Transactions are sequentially executed. Also, a transaction is either completed or reverted. Reversion means that the program state remains unchanged, but the attached gas is consumed.
Figure 2. Overview of the Ethereum system.
To help a user foresee a possible reversion and estimate the gas amount, the Ethereum system provides some free pre-execution services, such as infura.io (infura, 12). It is common for a TX issuer to send a transaction to such a service via its API, receive the pre-execution outcome, and let the user confirm if the transaction should be sent to the Ethereum network for real. The service deploys a node in the network, but no pre-execution result is actually committed.
2.2. Natures of smart contract computations
Smart contracts are mainly used for record-keeping, trading and arbitration purposes. Although the $GP$ is gigantic, the computation in each transaction has to be made much simpler than a typical computation in a traditional program, because it is orders-of-magnitude slower and more costly. In particular, the complexity due to loops and branches is much more manageable – Section 7.1 presents data on hundreds of real-world cases covering a substantial diversity in transactions between years 2022 and 2024. In nearly all these cases, loops are used to bundle a set of otherwise individual transactions into a batched transaction (for gas-saving and transactional guarantee). The majority of branches only cause a case split that is easy to verify jointly as a single theorem.
The main source of complexity, however, comes from the fact that cross-contract calls are made by object references, i.e. addresses. Essentially, the semantics of a contract code containing such a call is undefined until the address value is concretized at runtime.
///@custom:invariant sum(this.balances)==this.totalSupply
contract ERC20 {
address public owner;
uint256 public totalSupply
mapping(address => uint256) balances;
Declarations of transfer(...), allowance(...), transferFrom(...), ...
Code 1: Specifying Property 1 as the invariants for ERC20.
2.3. The NatSpec notation in Solidity
TCT uses the NatSpec notation of Solidity (NatSpec, 31) to specify interface specifications. A contract invariant is specified using the tag @custom:invariant above the contract’s definition, and a method postcondition is specified using the tag @custom:postcondition above the method’s definition. For example, LABEL:code:inv specifies Property 1 as invariants for the ERC20 contract. The Solidity compiler exports these contents as meta-data for the compiled bytecode.
3. Motivating Example
Figure 3. Attacks via integer overflow and reentrancy.
@@ 3. Motivating Example
To motivate our work, this section describes a token contract that combines two vulnerabilities – the first is exploitable via integer overflow, similar to the one described in reference (peckshield-blog, 23); the second is exploitable via reentrancy, similar to the DAO attack (reentrancy_attacks, 25). In Figure 3, the middle portion shows the contract code MultiVulnToken. The require statement checks a condition: if it is false, the execution is reverted. The Solidity keyword for method is function, so we use the terms “method” and “function” interchangeably.
3.1. Vulnerabilities and attacks
@@
In Attack 1, the attacker issues a transaction by calling transferProxy with arguments $value=2^{255}+1$ and $fee=2^{255}$ (in transition A). In the line require(balances[_from] >= _fee + _value), the two values are added, which equals 1 due to integer overflow, so the require statement is satisfied. Address msg.sender is the attacker’s address, and address _to is specified by the attacker. When the function returns, the balances of the two addresses become enormous (as transition B). This causes extreme inflation, so the attacker gets almost the entire asset of this token ecosystem.
@@
Attack 2 results in the attacker receiving 50 tokens in one of his accounts by clearing his other account that has only 5 tokens. It is achieved by making reentrant calls 10 times. The steps are as follows. First, the attacker creates a contract to call function clear in MultiVulnToken. Inside the function, msg.sender refers to the attacker’s contract. Therefore, the function call msg.sender.receive- Notification (transition C) calls into the attacker’s contract, who calls function clear again (transition D). This unexpected code sequence causes balances[_to]+=bal to be executed 10 times.
@@
A dilemma for static verification. Consider the perspective of a static verifier at compile time. The behavior of clear is not fully defined because it calls into another contract using an address (i.emsg.sender). The verifier can perhaps err by conservatively flagging all such calls as “potentially unsafe”. However, calling a contract via an address is a norm in smart contracts. Imagine another call businessPartner.foo(), where businessPartner is a variable only writable by the contract owner. Treating it as an arbitrary address, thus flagging the code as “potentially unsafe”, is problematic because all cross-contract calls would similarly be flagged “potentially unsafe”, which are false alarms. TCT is not subject to this dilemma because it is faithful to the $GP$ at runtime.
3.2. Theorem to prove safety of a transaction